OpenStudy (anonymous):

The square of the sum of two positive consecutive integers is 25. Find the two integers

OpenStudy (anonymous):

let first no is x then second would be x+1

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

what should be next condition ?

OpenStudy (anonymous):

according to your question

OpenStudy (anonymous):

0=x^2+x-25

OpenStudy (anonymous):

yes go on solve it by factorization or by formula

OpenStudy (anonymous):

is it 5 and -5

OpenStudy (anonymous):

sorry you made wrong equation

OpenStudy (anonymous):

ok now I know one of the answers is 5 but what is the other because its asking for two positive cosecutives

OpenStudy (anonymous):

so what is the right equation

OpenStudy (anonymous):

please check your procedure ... it is not correct .

OpenStudy (anonymous):

x^2 + x - 6 = 0

OpenStudy (anonymous):

please see here , first number = x second number be x + 1 so by the condition x^2 + (x+1)^2 = 25

OpenStudy (anonymous):

so you get ans 2 or -3

OpenStudy (anonymous):

first simp0lify the parentheses... we get x^2 + x^2 + 2x + 1 = 25

OpenStudy (anonymous):

@Shatish is doing write

OpenStudy (anonymous):

2x^2 + 2x = 24 x^2 + x = 12

OpenStudy (anonymous):

x^2 + x - 12 = 0 solve it and find the value of x .

OpenStudy (anonymous):

ok I know what i did wrong.. I always forget theres a ^2

OpenStudy (anonymous):

you will get two numbers are as 3 , 4

OpenStudy (anonymous):

i got 2 and -3 @Shatish

OpenStudy (anonymous):

are you getting it or not ?

OpenStudy (anonymous):

please check it again

OpenStudy (anonymous):

yes I get it. Thank you

OpenStudy (anonymous):

most welcome :)

OpenStudy (anonymous):

go and give medal to that person you think is able for that means give to @Shatish